A ball of mass 0.2 kg is thrown vertically upwards by applying a force by hand.If hand moves 0.2 m while applying the force and ball goes up to 2 m height further find magnitude of force. (g=10m/s^2)

Grade:11

1 Answers

Arun
6 years ago

We will solve it by energy conservation….

Work done by hand will change into potential energy=mgh

So work = f*s=mgh

Here s would be 0.2 meter for force by hand.. (force is applied while ball is in contact with hand)

And gravity will still work while ball is in contact with hand so h would be 2+0.2 meter

Thus f*0.2 =0.2*10*2.2( i have taken g =10m/s²)

F =22 N